OpenStudy (anonymous):

Solve, showing all steps: 3 - 5x > -8

OpenStudy (anonymous):

i need help please i need a awnser

OpenStudy (igreen):

You will subtract three..not add 8. @vzfreakz

OpenStudy (anonymous):

5/5=-5x/5

OpenStudy (anonymous):

@iGreen is right start from the opposite of the x

OpenStudy (anonymous):

you would want to leave x alone

OpenStudy (igreen):

\(3 - 5x > -8\) \(-5x > -8 - 3\) \(-5x > -11\) \(x < \dfrac{-11}{-5}\) or \(x < 2.2\) Remember to switch the sign when dividing by a negative number.

OpenStudy (anonymous):

can you give me the awnser please i need this i green

OpenStudy (igreen):

I did..

OpenStudy (anonymous):

@iGreen wrong X would equal to X<11/5 . There is no negative

OpenStudy (igreen):

There can be a negative. 11/5 and -11/-5 are the same thing.

OpenStudy (igreen):

\(\dfrac{11}{5}\) and \(\dfrac{-11}{-5} = 2.2\)

OpenStudy (anonymous):

well it would confuse people if you add negative for the nominator and denominator. Your right but the simpler the better ;-D

OpenStudy (igreen):

Yep, I agree. @TwinYang

OpenStudy (anonymous):

so it would be

OpenStudy (anonymous):

like @iGreen 11/5 or −11/−5 either one is the right answer.

OpenStudy (igreen):

Or you can say 2.2.

OpenStudy (anonymous):

lol yeah 11/5 , -11/-5 or 2.2.